A heater with a power of 150W is left on for 5 minutes after the water boils. Calculate the mass of water, in unit gram, that has become vapour during the period of 5 minutes.
[Given L of steam = 2260000Jkg-1]
1
Expert's answer
2017-09-14T07:06:06-0400
The amount of energy that transferred from heater to water
E=Pt.
On the other hand
E=Lm.
Thus
m=Pt/L=(150×5×60)/2260000=0.012 kg=12 g.
Answer: 12 g